Does the following situation represent a linear or exponential model? Write a function to model the situation.
Cooper has $75 in savings and is adding $15 each week.
How much will he have in 16 weeks?
Linear or Exponential?
After 16 weeks

Answer by jorel1380(3719)

Cooper contribution to his savings is constant, so the situation can be represented by a linear model. Thus:
Savings=75 + 15t...where t is the time in weeks. So, after 16 weeks we have:
S=75+15(16)=$315 total savings. ☺☺☺☺